OpenStudy (anonymous):

A coin purse contains a mixture of 33 dimes and quarters. The coins have a total value of $6.45. For the following problems, Use x for the number of dimes and y for the number of quarters.

OpenStudy (anonymous):

A)Write a mathematical expression for the total number of coins. B)Write a mathematical expression for the total monetary value in the purse. C)How many dimes are in the coin purse?

OpenStudy (ranga):

No. of dimes + No. of quarters = 33 x + y = 33 The value of x dimes is: 0.10 * x dollars The value of y quarters is: 0.25 * y dollars Total value = .1x + .25y = 6.45 A) x + y = 33 B) .1x + .25y = 6.45 For C) use the equations in A and B, eliminate y and solve for x

OpenStudy (anonymous):

i get x is 12

OpenStudy (anonymous):

but that isent the right anwser